if soccer field is 400 feet wide and 600 feet long how many feet to go diagonally across?
Answer by colliefan(242)

Use the Pythagorean theorum: a^2 + b^2 = c^2
400^2 + 600^2 = c^2
160000 + 360000 = c^2
520000 = c^2
sqrt(520000) = c
c= 721.11 feet